OpenStudy (anonymous):

degree of 2 zeros: 1-2i

OpenStudy (anonymous):

much easier than it looks there is an easy way, and a really really easy way lets do the easy way first

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

work backwards, like solving a quadratic equation backwards \[x=1-2i\] subtract \(1\) \[x-1=2i\] square both sides (carefully) \[(x-1)^2=(2i)^2\\ x^2-2x+1=-4\] then add \(4\) and finish with \[x^2-2x+5\]

OpenStudy (anonymous):

want to try the really really easy way?

OpenStudy (anonymous):

sure !

OpenStudy (anonymous):

it is really really easy but it requires memorizing something if the zero of a quadratic is \(a+bi\) then the quadratic is \[x^2-2ax+(a^2+b^2)\] you had \(1+2i\) so the quadratic is \[x^2-2\times 1x+(1^2+2^2)=x^2-2x+5\]

OpenStudy (anonymous):

what about when it tell you .. degree 3 zeros -1 with a multiplicity of 3 ?

OpenStudy (anonymous):

is -1 a and 3 b ?